### v1

# Electromagnetic Flow Control in Hypersonic Rarefied Environment

Abstract
The Unified Gas-Kinetic Wave-Particle (UGKWP) method, developed for multiscale simulation of partially ionized plasmas, has been extended to unstructured meshes, enabling the modeling of electromagnetic flows around a hemisphere across near-continuum to rarefied regimes. To the best of our knowledge, this work represents the first application of a multiscale plasma solver to this problem. In our approach, neutrals, ions, and electrons are treated as distinct species, with electrons modeled beyond the conventional fluid approximation. The numerical implementation is validated through comparison with reference solutions for neutral hypersonic flow around a sphere, as well as benchmarking against experimental data for a Mach 4.75 pre-ionized argon flow. In both cases, the UGKWP results show good agreement with the reference and experimental data. The findings reveal that rarefied effects play a significant role in the prediction of electromagnetic flow control, underscoring the necessity of multiscale modeling in plasma flow applications.
keywords:
unified gas-kinetic wave-particle method , partially-ionized plasma , electromagnetic flow control , multiscale algorithm
journal: Elsevier
1 Introduction
Magnetohydrodynamic (MHD) thermal protection technology holds strategic significance for both national defense systems and deep-space exploration missions [1]. During hypersonic flight, a strong externally applied magnetic field interacts with the weakly ionized gas produced by intense shock waves, increasing the shock standoff distance and thereby reducing the heat flux to the vehicle surface. The theoretical foundations of this technology date back to the 1950s US-Soviet space race, when Resler and Sears pioneered the formulation of fundamental magneto-aerodynamic equations [2]. They modeled weakly ionized plasma in high-speed, high-enthalpy flows as conducting fluids, incorporating Lorentz force and Joule heating terms into the modified Navier–Stokes equations.
Over subsequent decades, considerable progress was made in theoretical, computational, and experimental aspects of MHD thermal protection. In 1958, Bush provided an analytical solution describing the magnetic field’s effect on shock standoff distance [3], while Ziemer and Bush later provided experimental validation using a Mach 4.5 sphere model [4], marking important early milestones. Cambel further investigated the Hall effect’s detrimental impact on magnetic control efficiency [5] and conducted corresponding experiments using pre-ionized argon gas [6], creating benchmark datasets that remain valuable for numerical validation. Following these foundational studies, advances in computational fluid dynamics enabled high-fidelity simulations of MHD flows [7, 8, 9]. Recent numerical studies have increasingly incorporated complex physical phenomena, such as rarefied gas effects and wall sheath interactions. For example, Fawley [10] utilized Direct Simulation Monte Carlo (DSMC) methods to model conductivity in transitional flow regimes, while Katsurayama [11] applied DSMC to simulate rarefied gas effects, approximating the Lorentz force as a body force rather than explicitly resolving charged particle trajectories. Parent [12] further quantified the impact of plasma sheaths on the overall flow field.
Despite these advancements, most numerical simulations rely on MHD equations, whose applicability diminishes under high-altitude conditions characterized by large Knudsen numbers. In such regimes, the continuum assumption inherent to MHD-based models breaks down [13]. Accurately capturing flow physics across continuum, transitional, and rarefied regimes thus requires a genuinely multiscale numerical approach. Addressing these challenges, recent developments have led to the Unified Gas-Kinetic Wave-Particle (UGKWP) method, which enables robust modeling of multiscale neutral gas flows from the rarefied to continuum regimes [14, 15]. The UGKWP framework has since been systematically extended to a variety of transport phenomena, including plasma [16, 17], granular flows [18, 19], radiative transfer [20, 21], and phonon transport [22], demonstrating its capability to capture non-equilibrium effects. Most recently, UGKWP has been specifically applied to partially ionized plasma flows [23]. In the continuum regime, UGKWP accurately recovers ideal MHD, dissipative MHD, and two-fluid models; as the flow transitions to rarefied conditions, the method seamlessly shifts to a particle-based description, effectively capturing kinetic plasma behavior and recovering Particle-in-Cell (PIC) dynamics in the collisionless limit.
The remainder of this paper is organized as follows: Section 2 provides a concise overview of the UGKWP method for partially ionized plasma. Section 3 presents the numerical test results. Finally, Section 4 offers the main conclusions of this study.
2 UGKWP Method
The UGKWP method is used to perform numerical simulations in this study. In the framework of finite volume method, the cell-averaged macroscopic conservative variable $\boldsymbol{W}_{\alpha}=(\rho_{\alpha},\rho_{\alpha}\boldsymbol{U}_{\alpha},\rho_{\alpha}\mathscr{E}_{\alpha})$, i.e, mass, momentum and energy densities of species $\alpha$ on a physical cell $\Omega_{i}$ are updated by the flux through cell interfaces, electromagnetic forces and cross-species collisions:
$(\boldsymbol{W}_{\alpha})_{i}^{n+1}=(\boldsymbol{W}_{\alpha})_{i}^{n}-\frac{\Delta t}{|\Omega_{i}|}\sum_{s\in\partial\Omega_{i}}|l_{s}|(\mathscr{F}_{\boldsymbol{W}_{\alpha}})_{s}+\frac{\Delta t}{\tau_{\alpha}}(\bar{(\boldsymbol{W}}_{\alpha})_{i}^{n}-(\boldsymbol{W}_{\alpha})_{i}^{n})+\Delta t(\boldsymbol{S}_{\alpha})_{i}^{n+1},$ (1)
where $l_{s}\in\partial\Omega_{i}$ is the cell interface with center $\boldsymbol{x}_{s}$ and outer unit normal vector $\boldsymbol{n}_{s}$. $|l_{s}|$ is the area of the cell interface. $(\boldsymbol{S}_{\alpha})_{i}$ is source term due to electromagnetic force. The numerical flux across interface $(\mathscr{F}_{\boldsymbol{W}_{\alpha}})_{s}$ can be evaluated from distribution function at the interface,
$(\mathscr{F}_{\boldsymbol{W}_{\alpha}})_{s}=\frac{1}{\Delta t}\int_{t^{n}}^{t^{n+1}}\boldsymbol{u}\cdot\boldsymbol{n}_{s}f_{\alpha}(\boldsymbol{x}_{s},\boldsymbol{u},\boldsymbol{\xi},t)\boldsymbol{\Psi}\mathrm{d}\boldsymbol{\Xi}\mathrm{d}t,$ (2)
where $\boldsymbol{\Psi}=(1,\boldsymbol{u},\frac{1}{2}(\boldsymbol{u}^{2}+\boldsymbol{\xi}^{2}))$ is the conservative moments of distribution functions with $\boldsymbol{\xi}=(\xi_{1},\xi_{2},\cdots,\xi_{n})$ the internal degree of freedom. $\mathrm{d}\boldsymbol{\Xi}=\mathrm{d}\boldsymbol{u}d\boldsymbol{\xi}$ is the volume element in the phase space. For cell interface $l_{s}$, the distribution function $f_{\alpha}(\boldsymbol{x}_{s},\boldsymbol{u},\boldsymbol{\xi},t)$ cannot be analytically represented in the transitional and collisionless regimes, thus requiring direct tracking of its evolution. The evolution of the distribution function in non-continuum regimes is crucial for realizing multiscale transport in the UGKWP method, as detailed in [23]. $(\bar{\boldsymbol{W}}_{\alpha})_{i}=((\rho_{\alpha})_{i},(\rho_{\alpha}\bar{\boldsymbol{U}}_{\alpha})_{i},(\rho_{\alpha}\bar{\mathscr{E}}_{\alpha})_{i})$ where $(\bar{\boldsymbol{U}}_{\alpha})_{i}$ and (¯$(\bar{\mathscr{E}}_{\alpha})_{i}$ are post-collision velocity and energy in multispecies kinetic model. In this paper, the Andries-Aoki-Perthame (AAP) model is used to model cross-species collisions. In the AAP model, post-collision temperature and velocity are calculated as:
$\bar{\boldsymbol{U}}_{\alpha}$ $=\boldsymbol{U}_{\alpha}+\frac{\tau_{\alpha}}{m_{\alpha}}\sum_{k=1}^{N}2\mu_{\alpha k}\chi_{\alpha k}n_{k}\left(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha}\right),$ (3)
$\bar{T}_{\alpha}$ $=T_{\alpha}-\frac{m_{\alpha}}{3k_{B}}\left(\boldsymbol{U}_{\alpha}^{*}-\boldsymbol{U}_{\alpha}\right)^{2}+\tau_{\alpha}\sum_{k=1}^{N}\frac{4\mu_{\alpha k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}}\left(T_{k}-T_{\alpha}+\frac{m_{k}}{3k_{B}}\left(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha}\right)^{2}\right),$
where $\mu_{\alpha k}=m_{\alpha}m_{k}/(m_{\alpha}+m_{k})$ is reduced mass, mean relaxation time $\tau_{\alpha}$ is determined by $1/\tau_{\alpha}=\sum_{k=1}^{m}\chi_{\alpha k}n_{k}$, and interaction coefficient $\chi_{\alpha k}$ for hard sphere model is [24]:
$\chi_{\alpha k}=\frac{4\sqrt{\pi}}{3}\left(\frac{2k_{B}T_{\alpha}}{m_{\alpha}}+\frac{2k_{B}T_{k}}{m_{k}}\right)^{1/2}\left(\frac{d_{\alpha}+d_{k}}{2}\right)^{2},$
where $d_{\alpha},d_{k}$ are the diameters of the particles and can be approximated by
$(d_{\alpha}+d_{k})^{2}=\frac{1}{\sqrt{2}\pi(n_{\alpha}+n_{k})\text{Kn}\text{L}},$
where Kn is Knudsen number, L is reference length. The molecular diameter is modified by the temperature as
$\left(\frac{d}{d_{ref}}\right)=\left(\frac{T}{T_{ref}}\right)^{-\xi/2},$
where $\xi=\omega-\frac{1}{2}$, and $\omega$ is the power coefficient between viscosity and temperature[25]. In this work, $\omega=0.81$ for argon atom. The cross section between electron and neutral particles is described by a fitted curve from Devoto [26].
In this study, the low magnetic Reynolds number approximation is employed, which implies that the externally applied magnetic field dominates over the induced magnetic field generated by the current. This assumption holds when the electrical conductivity is sufficiently low. Additionally, an approximate Debye length is implemented to alleviate numerical stiffness, as resolving the exact Debye length would be computationally prohibitive, particularly in three-dimensional simulations.
3 Numerical Studies
3.1 Hypersonic Flow Around a Sphere
[…]
Parameter Value
Mach number 4.25
Knudsen number 0.031
Argon atomic mass $6.63\times 10^{-26}\ \mathrm{kg}$
Argon diameter (VHS) $4.17\times 10^{-10}\ \mathrm{m}$
Incoming argon density $1.3841\times 10^{-6}\ \mathrm{kg/m^{3}}$
Characteristic temperature 273 K
[…]
Wall temperature 273 K
Incoming argon velocity 1308.21 m/s
Sphere diameter 2 m
Prandtl number 1
Table 1: Simulation parameters for the rarefied argon flow around a sphere.
Figure 1 displays the contour distributions of density, velocity, temperature, and local Knudsen number on and around the wall. The significant variation in Knudsen number, from 0.031 to values greater than 1 across the entire flow field (Figure 1(d)), highlights the algorithm’s effectiveness in capturing flow physics across different rarefied flow regimes. The surface quantities are presented in Figure 2. The present results exhibit excellent agreement with the reference data generated by the Unified Gas-Kinetic Scheme (UGKS) code[27]. This shows the accuracy of the current solver in the ionization-free condition.
Figure 1: Hypersonic flow around a sphere at Ma=4.25 and Kn=0.031 by the UGKWP method. a) density, b) velocity in x direction, c) temperature, and d) local Knudsen number contours.
(a) Surface Pressure Coefficient
(b) Surface Heat Flux Coefficient
(c) Surface Shear Stress Coefficient
Figure 2: Distributions of Surface Pressure, Heat Flux, and Shear Stress, the reference data is produced by the UGKS code.
3.2 Hypersonic Pre-ionized Argon Flow Around a Hemisphere
To further assess the code’s accuracy, we conduct a three-dimensional simulation of Mach 4.75 argon flow around a hemisphere, replicating the experimental conditions reported by Kranc et al. [6]. The hemisphere radius is defined as $r_{n}=0.01905$ m. Figure 3(a) illustrates the computational domain geometry employed in this study. The simulation imposes a hypersonic freestream flow in the $-x$ direction incident upon the hemisphere. We generate an unstructured hexahedral mesh to discretize the computational domain surrounding the hemisphere.
(a) Computational domain geometry
(b) Dipole magnetic field contours ($B_{\text{max}}=0.387\ \text{T}$).
Figure 3: Computational domain specifications and magnetic magnitude distribution for hypersonic flow over a hemisphere.
In the experiments conducted by Kranc et al., the inflow comprised a partially ionized argon flow generated by a plasma torch. Kranc et al. estimated the ionization degree to be $\alpha=0.025$, and then Bisek et al. [28], employing a more refined Saha equation for weakly ionized atomic gases, re-estimated the ionization degree as $\alpha=0.00623$. This revised value is adopted in the present study. Kranc et al. asserted that the flow remained chemically frozen, allowing the ionization degree to be treated as a constant throughout the flow field. Detailed parameters are listed in Table 2.
Parameter Value
Mach number 4.75
Knudsen number 0.044
Argon atomic mass $6.63\times 10^{-26}\ \mathrm{kg}$
Argon diameter (VHS) $4.17\times 10^{-10}\ \mathrm{m}$
Incoming argon density $1.09\times 10^{-4}\ \mathrm{kg/m^{3}}$
Incoming argon ion density $6.85\times 10^{-7}\ \mathrm{kg/m^{3}}$
Incoming electron number density $1.0332\times 10^{19}\ \mathrm{m^{-3}}$
Characteristic temperature 1100 K
Incoming electron temperature 5000 K
Wall temperature 1100 K
Incoming argon velocity 3000 m/s
Hemisphere radius 0.01905 m
Prandtl number 1
Table 2: Simulation parameters for the rarefied argon flow around a hemisphere.
A magnetic dipole, with its magnetic moment oriented along the x-axis, is positioned at the center of the hemisphere. Its orientation opposes the incoming flow along the stagnation line. The magnetic field generated by this dipole is defined as:
$\mathbf{B}=\frac{B_{\max}}{2\left(\hat{x}^{2}+\hat{y}^{2}+\hat{z}^{2}\right)^{5/2}}\left[\begin{array}[]{c}2\hat{x}^{2}-\left(\hat{y}^{2}+\hat{z}^{2}\right)\\ 3\hat{x}\hat{y}\\ 3\hat{x}\hat{z}\end{array}\right],$
where, $\hat{x}=x/r_{n}$, $\hat{y}=y/r_{n}$, and $\hat{z}=z/r_{n}$ represent the coordinates normalized by the hemisphere radius, $r_{n}$. The parameter $B_{max}$ denotes the maximum magnetic field strength, which is attained at the stagnation point. The magnetic field magnitude distribution is depicted in Figure 3(b). In the following analysis, we employ $B_{max}$ to characterize the magnetic field strength.
Simulations are performed at four distinct magnetic field strengths: 0.223 T, 0.316 T, 0.387 T, and 0.447 T. The simulations employ a two-step approach. Initially, the flow field is computed in the absence of a magnetic field until a steady-state solution is achieved. Subsequently, the magnetic field is introduced, and the simulation is continued until a new steady-state solution, reflecting the influence of the applied magnetic field, is obtained.
Figure 4: Contour plots of (a) density, (b) x-velocity, (x) temperature, and (d) heat flux, comparing conditions with and without an applied magnetic field. The upper half of each plot shows results for $B_{0}=0T$, while the lower half illustrates the effects of a $B_{0}=0.387T$ background magnetic field. The data clearly demonstrate that the magnetic field increases the shock standoff distance and markedly decreases the heat flux.
Figure 5: Contour plots on the wall of heat flux of (a) Ar, (b) Ar+, (c) electron, comparing conditions with and without an applied magnetic field. The upper half of each plot shows results for $B_{0}=0T$, while the lower half illustrates the effects of a $B_{0}=0.387T$ background magnetic field. The data clearly demonstrate that the magnetic field decreases the heat flux of Ar, but increases that of Ar+ and electrons.
Kranc et al. [6] observed an increase in shock standoff distance when an external magnetic field was applied. This increase was determined by comparing photographic images of the flow with and without the magnetic field, with the upstream edge of the shock inferred from the boundary of the flow’s luminosity. In computational simulations, a universally accepted definition of the shock edge remains elusive. Bisek et al. [28] defined the shock location as the point along the stagnation line where the density ratio exceeds the theoretical infinite Mach number threshold for a normal shock wave in an ideal gas. However, this density ratio proves excessively high for rarefied flow conditions. In the present study, we estimate the shock edge location to occur where the density ratio exceeds twice the freestream density, a criterion that aligns with the shock edge observed in density contours.
Contour distributions of density, velocity, temperature, and heat flux on and around the wall are presented in Figures 4 and 5. The upper half of each plot represents the results without a magnetic field ($B_{0}=0\text{ T}$), and the lower half shows the results of a $B_{0}=0.387\text{ T}$ background magnetic field. The magnetic field clearly increases the shock standoff distance. For argon, a smaller temperature gradient leads to a reduction in heat flux. Conversely, the heat flux for argon ions ($\text{Ar}^{+}$) and electrons ($\text{e}^{-}$) increases, possibly due to direct magnetic heating. However, the overall heat flux is lower due to the low concentration of $\text{Ar}^{+}$ and $\text{e}^{-}$ relative to Ar, and the minor increment in heat flux is trivial compared to the decrement observed in Ar.
Figure 6 depicts the variation in shock standoff distance with changing argon ion-electron mass ratios in our simulations. We observe a gradual convergence of the standoff distance as the mass ratio increases. Critically, our calculations, performed with the multiscale solver UGKWP to accurately capture the Ar flow field, show better agreement with the experimental results of Kranc et al. This contrasts with reference computational results from MHD employing different conductivity models. The findings clearly highlights that an accurate Ar flow field is fundamental to accurately predicting the shock standoff distance.
Figure 6: Changes in shock standoff distance as a function of magnetic field strength for Mach 4.75 pre-ionized argon flow around a hemispherical geometry, parameterized by mass ratio. Experimental data are from Kranc et al. [6], while computational results using MHD methods are from Bisek et al. [28] using various conductivity model. (Experimental uncertainties is ±$\pm 10\%$)
4 Conclusions
In this study, the UGKWP method has been employed for the multiscale simulation of partially ionized plasma flows around a hemisphere, spanning regimes from near-continuum to highly rarefied. Validation against established UGKS codes for neutral flows, as well as experimental data for pre-ionized argon flows, demonstrates the accuracy and reliability of our numerical implementation. The results reveal that rarefied effects have a substantial impact on the prediction of electromagnetic flow control, underscoring the necessity of multiscale modeling in plasma flow applications.
Future work will focus on incorporating additional physical phenomena, such as ambipolar diffusion, shock polarization, wall sheath effects arising from small Debye lengths, and comprehensive magnetohydrodynamic effects for moderate to large magnetic Reynolds numbers. The fidelity of these advanced simulations will continue to depend on the robust flow field solutions provided by the UGKWP multiscale solver. Overall, this research establishes a solid foundation for the development of more sophisticated plasma flow simulations in the future.
5 Acknowledgement
This current work was supported by National Key R&\&&D Program of China (Grant Nos. 2022YFA1004500), National Natural Science Foundation of China (12172316, 92371107), and Hong Kong research grant council (16301222, 16208324).